# ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA

### v1

@@ Abstract
Recent advances in large language models (LLMs) have significantly improved multi-hop question answering (QA) through direct Chain-of-Thought (CoT) reasoning. However, the irreversible nature of CoT leads to error accumulation, making it challenging to correct mistakes in multi-hop reasoning. This paper introduces ReAgent: a Reversible multi-Agent collaborative framework augmented with explicit backtracking mechanisms, enabling reversible multi-hop reasoning. By incorporating text-based retrieval, information aggregation and validation, our system can detect and correct errors mid-reasoning, leading to more robust and interpretable QA outcomes. The framework and experiments serve as a foundation for future work on error-tolerant QA systems. Empirical evaluations across three benchmarks indicate ReAgent’s efficacy, yielding average about 6% improvements against baseline models.
ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA
Zhao Xinjie1, Fan Gao1, Rui Yang2, Yingjian Chen1, Yuyang Wang3, Ying Zhu1, Jiacheng Tang1, Irene Li1 1 The University of Tokyo, 2 Duke-NUS Medical School, 3 The Hong Kong University of Science and Technology (Guangzhou)
@@ 1 Introduction
Multi-hop question answering (QA) represents a crucial challenge in natural language processing (NLP), requiring systems to integrate and reconcile evidence from multiple sources or hops of reasoning before reaching a single final answer (Yang et al., 2018; Welbl et al., 2018; Ho et al., 2020). Existing benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020) have shed light on the intrinsic difficulty of multi-hop inference: not only must relevant information be located across dispersed documents or knowledge graphs, but the process itself must remain robust to partial mistakes. If an early inference is incorrect, subsequent steps can amplify that error and produce an unreliable or contradictory final answer (Inoue et al., 2020; Bo et al., 2024; Yang et al., 2024c).
Recent large language models (LLMs) have showcased strong performance on multi-hop QA, often employing either explicit Chain-of-Thought (CoT) prompting (Wang et al., 2023) or retrieval-augmented generation (Das et al., 2018; Long et al., 2019; Gao et al., 2023; Liu et al., 2023; Yang et al., 2024b, 2025). While these strategies can help track intermediate steps, they typically rely on a predominantly forward-directed workflow. Once a partial conclusion is formed, the system rarely revisits or corrects it unless externally prompted (Puerto et al., 2023; Huang et al., 2024). This forward-only structure poses limitations, especially when later evidence contradicts an earlier inference or when multiple lines of reasoning yield divergent findings. A single incorrect hop can invalidate the final answer, and without explicit backtracking, the system is unable to retract and repair flawed assumptions (Doyle, 1979; Bo et al., 2024).
Figure 1: We compare different frameworks for QA: Linear (CoT), Collaborative (MAS), and Reversible (ReAgent) Reasoning in Multi-hop QA.
To address this gap, growing attention has been directed to multi-agent collaboration for QA, where specialized agents focus on sub-tasks—such as retrieval, verification, and resolution of conflicting information—and then synchronize their outcomes (Zhao et al., 2024; Parhizkar et al., 2020; Ke et al., 2024). By delegating different roles (e.g., question decomposition, result verification) to different agents, the system becomes more modular and potentially more explainable. Recent multi-agent systems include COPPER (Bo et al., 2024) and LongAgent (Zhao et al., 2024), both of which decompose complex tasks among multiple LLM-based agents that coordinate via message passing. They not only reduce overall context per agent but can also cross-check for consistency, flagging suspicious partial solutions. However, even these advanced frameworks often lack fully reversible reasoning pipelines, wherein each agent can effectively backtrack to an earlier valid state once a conflict is detected, re-explore alternatives, and propagate the corrections forward (He et al., 2021).
In this paper, we introduce ReAgent, a reversible multi-agent collaborative reasoning framework with explicit backtracking, designed for multi-hop question answering. This approach builds on the premise that in real-world complex QA scenarios—especially when reasoning across a broad knowledge graph or multiple textual sources—one-size-fits-all steps fail to capture the frequent need for partial re-evaluation (Yang et al., 2018; Ho et al., 2020). Conflict detection and backtracking mechanisms act as a flexible error-correction loop, in which a mismatch in evidence can trigger a re-check of previously accepted facts, ensuring that misinterpretations do not survive to the final answer. This perspective aligns with “reflective” multi-agent designs (Bo et al., 2024; Huang et al., 2024), but pushes further by defining explicit multi-layer protocols—local versus global backtracking—and employing a dynamic aggregator that leverages trust signals from each agent’s past reliability (Puerto et al., 2023; Parhizkar et al., 2020).
We empirically show that this reversible approach improves final-answer accuracy and interpretability: backtracking traces clarify how the system isolates and corrects mistakes. Our work links robust non-monotonic reasoning with data-driven QA, offering a transparent, error-tolerant multi-hop inference pipeline. In summary, our main contributions are: (1) a novel multi-agent QA framework that supports backtracking at both local and global levels; (2) hybrid retrieval from knowledge graphs and text for flexible evidence acquisition; and (3) systematic experiments on three benchmark datasets demonstrating more robust and interpretable multi-hop QA than forward-only baselines. Specifically, our framework achieves an average improvement of approximately 6% over baseline models.
Figure 2: Execution, Supervision, and Interaction Layers: Three-Fold Role Partitioning in ReAgent’s Multi-Hop QA Backtracking Mechanisms.
@@ 2 Preliminary
Challenges in Forward-Only Reasoning A key source of error in multi-hop QA lies in the lack of revision capacity. Once a system commits to a specific inference early on (for instance, deducing a partial conclusion from $e_{1}$ and $e_{2}$), there is no built-in mechanism to retract or update that inference if additional evidence suggests it is incorrect. As the system progresses to hop $j{+}1$ or further, it carries potential inaccuracies along, making the final answer unreliable or inconsistent. In many cases, the model would have benefited from re-examining initial assumptions, especially when encountering contradictory or low-confidence signals. These observations motivate an interactive and flexible approach, wherein the QA process is allowed to backtrack and replace problematic intermediate results.
Non-Monotonic Backtracking Our notion of backtracking is a non-monotonic extension of the typical multi-hop QA process. Rather than treating the knowledge state as strictly accumulative, we allow for partial or global reversion. Specifically:
Local Correction: An agent can retract one of its own inferences if it detects internal conflicts or if it receives contradictory evidence from other agents.
Global Rollback: When inconsistencies span multiple agents, a supervisory mechanism coordinates a synchronized rollback. This can mean discarding a sequence of accepted statements that are ultimately untenable, returning the system to a previously consistent state.
This extended capability is guided by two key considerations. First, not all new evidence is guaranteed to be more credible than old; hence, the system must weigh or prioritize contradictory claims. Second, if an inference $I$ made in hop $r$ is discredited in hop $r{+}2$, it may be more efficient to rollback to hop $r$ directly, rather than forcing the pipeline to proceed with a known contradiction. Such rollback actions effectively prevent one faulty hop from invalidating the entire chain of reasoning.
Figure 3: Layered Multi-Agent Interaction and Backtracking Flow with Time Step.
@@ 3 ReAgent: Reversible Multi-Agent Reasoning Architecture
3.1 Overview of the ReAgent Architecture
Figure 2 presents the overall design, organized into three layers. Let $\mathcal{A}\;=\;\{A_{1},A_{2},\dots,A_{n}\}$ be the set of agents. At time step $t$, each agent $A_{i}$ maintains a local knowledge set $\Phi_{i}^{t}$ that may include proposed assertions, retrieved evidence, or partial inferences. In addition, the system can track a global assertion set $\Phi^{t}\;=\;\bigcup_{i=1}^{n}\Phi_{i}^{t},$ representing all statements under consideration at time $t$. The three-layer structure is as follows:
  1. 1.
    Execution Layer (Section 3.2): Comprises the essential agents for domain tasks, such as decomposing questions, retrieving evidence, verifying local consistency, and assembling intermediate answers.
  2. 2.
    Supervisory Layer (Section 3.3): Includes supervisory agents that orchestrate conflict resolution and manage global backtracking, ensuring that contradictory or invalid assertions are addressed across all relevant agents.
  3. 3.
    Interaction Layer (Section 3.4): Maintains the concurrency model and communication protocols. It defines how agents send and receive messages (e.g., assert, reject, challenge) and how the system merges these parallel events into consistent updates of $\Phi^{t}$.
By separating concerns across these layers, it becomes simpler to reason about local agent updates while still allowing global interventions. The design explicitly supports non-monotonic updates: newly introduced statements can be revoked if they lead to logical conflicts or are superseded by contradictory evidence. The specific prompts of each kind of agent is provided in Appendix A.
@@ 3.2 Execution Layer Agents
The Execution Layer hosts four types of agents that address fundamental QA subtasks, each maintaining its local knowledge $\Phi_{i}$. When necessary, an agent may perform local backtracking to remove or adjust recent assertions.
Question-Decomposer Agent ($A_{Q}$).
Breaks a complex input question $Q_{0}$ into sub-questions $\{q_{1},q_{2},\dots\}$. If $Q_{0}(x)$ denotes the main query, $A_{Q}$ produces a decomposition mapping $\{q_{j}\}$ for subsequent retrieval and verification: $A_{Q}:\;Q_{0}(x)\;\longmapsto\;\{\,q_{1},\dots,q_{m}\}.$ This decomposition is broadcast to other base-layer agents.
Retriever Agent ($A_{R}$).
Upon receiving a sub-question $q_{i}$, $A_{R}$ returns relevant evidence $E$ (e.g., text passages or knowledge graph triples), possibly with confidence scores $\alpha(E)$. This agent may maintain a local retrieval cache and re-invoke searches if backtracking invalidates prior evidence.
Verifier Agent ($A_{V}$).
Performs local consistency checks. Given new evidence $E$, it verifies coherence with $\Phi_{V}^{t}$ (the Verifier’s current knowledge). If conflicts arise, $A_{V}$ triggers local backtracking to revert to a consistent state or signals higher-level agents for broader resolution. Concretely, $A_{V}$ may invoke
$\mathrm{BacktrackLocal}\bigl{(}A_{V},r\bigr{)}$
if it suspects that assertions introduced after node $r$ are inconsistent.
Answer-Assembler Agent ($A_{A}$).
Gathers partial answers (and verified evidence) to synthesize a final solution. If the set of local inferences from $A_{Q}$, $A_{R}$, and $A_{V}$ is $\Phi_{\text{base}}$, then $A_{A}$ combines them to generate the final answer $A_{\text{final}}$. Any detected contradiction triggers further checks or possible escalation to the supervisory layer.
@@ 3.3 Supervisory Layer Agents
The supervisory layer manages conflicts that cannot be resolved by a single agent. It also oversees system-wide strategies, especially when contradictory goals or inconsistent states appear across multiple agents.
Supervisor Agent ($A_{S}$).
Coordinates multi-agent conflicts spanning multiple knowledge sets or critical shared assumptions. If a conflict persists after local backtracking, $A_{S}$ determines whether partial or holistic rollback is required. When a system-wide contradiction is found, one may identify a minimal conflict set $\Psi$ such that $\mathrm{SAT}(\Psi)=\textsf{false}$ but $\mathrm{SAT}(\Psi^{\prime})=\textsf{true}$ for any proper subset $\Psi^{\prime}\subset\Psi$. The supervisor then triggers a backtracking operation that eliminates or modifies $\Psi$.
Controller Agent ($A_{C}$).
Regulates high-level strategies, monitoring game-theoretic signals or meta-rules. If certain rules appear suboptimal or unsafe, $A_{C}$ can override them or enforce mode switches. For instance, it may issue a $\textsf{challenge}(\varphi)$ to examine a crucial assertion $\varphi$ from multiple agents’ perspectives, or override local decisions if they jeopardize overall consistency.
@@ 3.4 Interaction Layer
Communication Protocols.
We adopt a message-driven framework, where atomic events $\mathtt{msg}(\varphi)$ are labeled with types such as assert, inform, reject, or challenge. These updates can occur simultaneously, requiring a composite event model to merge parallel messages. A shared concurrency mechanism processes simultaneous actions, ensuring that if conflicts arise from concurrent assertions, they are escalated to the supervisory layer.
Temporal Operators.
The system can utilize operators like $\square\varphi$ to denote that $\varphi$ must hold in every future state, and $\lozenge\varphi$ to indicate that a proposition $\varphi$ might become true at some future point. These temporal constraints assist in specifying persistent axioms or potential triggers for backtracking conditions.
Algorithm 1 Multi-Agent Reversible Reasoning
1:$Q_{0}$: Main question
$A_{Q}.\mathrm{decompose}(Q_{0})\to\{q_{1},q_{2},\dots\}$
3:Broadcast $\mathrm{assert}(q_{i})$ to $A_{R},A_{V},A_{A}$
4:for each $q_{i}$ in $\{q_{1},q_{2},\dots\}$ concurrently do
5:     $E\leftarrow A_{R}.\mathrm{retrieve}(q_{i})$
6:     $A_{V}.\mathrm{verify}(E)$
7:     if $A_{V}$ detects conflict locally then
8:         $A_{V}.\mathrm{BacktrackLocal}(r_{V})$
9:         if conflict persists then
10:              raise Conflict to $A_{S}$
11:         end if
12:     end if
13:     $A_{A}.\mathrm{storePartialAnswer}(q_{i},E)$
14:end for
15:if Global conflict signaled then
16:     $A_{S}.\mathrm{HolisticUpdate}(\Pi_{j})$
17:     if $A_{C}$ intervention needed then
18:         $A_{C}.\mathrm{challenge}(\varphi)$ or override
19:     end if
20:end if
22:return Final Answer

### v2

@@ Abstract
Multi-hop question answering (QA) remains challenging, as solutions must reliably integrate and reconcile evidence from multiple sources without succumbing to error propagation. While large language models (LLMs) have achieved substantial improvements via chain-of-thought (CoT) prompting and retrieval-augmented generation, these methods typically adopt a forward-only workflow—early mistakes persist throughout inference, and contradictions discovered later cannot systematically trigger re-evaluation. To address this limitation, we present ReAgent, a reversible multi-agent reasoning framework. Specifically, ReAgent enables agents to backtrack to earlier valid states when conflicts arise, thereby isolating and rectifying flawed assumptions before they undermine subsequent reasoning. Our approach combines explicit local and global rollback protocols with modular role specialization, resulting in a flexible and error-tolerant pipeline. Empirical evaluation on three multi-hop QA benchmarks demonstrates consistent performance gains of approximately 6% over forward-only baselines, in addition to enhanced interpretability. These findings highlight the value of non-monotonic, backtracking-driven inference in complex QA scenarios and point to broader implications for multi-agent collaboration in knowledge-intensive tasks. [^1] Our anonymous code is available at https://anonymous.4open.science/r/ReAgent-9415.
ReAgent: Reversible Multi-Agent Reasoning for
Knowledge-Enhanced Multi-Hop QA
Xinjie Zhao1$\dagger$, Fan Gao1$\dagger$, Xingyu Song1, Yingjian Chen1, Rui Yang2, Yanran Fu1, Yuyang Wang3, Yusuke Iwasawa1, Yutaka Matsuo1, Irene Li1 1The University of Tokyo 2Duke–NUS Medical School 3The Hong Kong University of Science and Technology irene.li@weblab.t.u-tokyo.ac.jp
$\dagger$\$$\dagger$\$footnotetext: Equal contribution.
@@ 1 Introduction
Multi-hop question answering (QA) is a central challenge in natural language processing (NLP), demanding the ability to gather and integrate evidence across multiple documents, database entries, or knowledge-graph nodes before converging on a single correct answer (Yang et al., 2018; Welbl et al., 2018; Ho et al., 2020). Benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020) highlight the intricate nature of multi-hop inference, where each reasoning step can involve partial retrieval, validation, and synthesis of new information. A core difficulty lies in the system’s vulnerability to early mistakes: if an incorrect inference is made at an initial hop, subsequent steps often propagate this error and undermine the final outcome, rendering it contradictory or simply wrong (Inoue et al., 2020; Bo et al., 2024; Yang et al., 2024c). This phenomenon has motivated extensive research into chaining intermediate inferences and exploring ways to detect, isolate, or rectify problematic conclusions.
Figure 1: Comparison of multi-hop reasoning strategies. Chain-of-Thought (CoT) and Multi-Agent Systems (MAS) typically adopt a forward-driven reasoning pipeline without rollback mechanisms, which could generate the wrong answer due to error accumulation. In contrast, our proposed ReAgent introduces explicit backtracking mechanisms that enable the system to correct errors during reasoning, resulting in a more accurate and reliable answer.
Recent large language models (LLMs) exhibit promising results on multi-hop QA, frequently using either explicit Chain-of-Thought (CoT) prompting (Wang et al., 2023) (Figure 1, left) or retrieval-augmented generation (Das et al., 2018; Long et al., 2019; Gao et al., 2023; Liu et al., 2023; Yang et al., 2024b, 2025a). These methods facilitate a stepwise approach, encouraging transparency in how each intermediate fact is reached. Nonetheless, they typically rely on a forward-driven reasoning pipeline that does not proactively examine or correct previously accepted statements. In practice, once a model commits an erroneous partial inference may not reevaluate it unless given targeted prompts to do so (Puerto et al., 2023; Huang et al., 2024). This unidirectional paradigm is problematic when later-discovered evidence or reasoning paths contradict prior assumptions, as the system lacks a structured mechanism to revise and propagate corrections. Although incremental improvements have been proposed, the absence of robust backtracking or rollback still limits their reliability and interpretability (Doyle, 1979; Bo et al., 2024).
A growing body of work in multi-agent collaboration (Figure 1, middle) offers an alternative perspective, assigning specialized components distinct roles in retrieval, validation, conflict detection, and results assembly (Zhao et al., 2024; Parhizkar et al., 2020; Ke et al., 2024). Approaches such as COPPER (Bo et al., 2024) and LongAgent (Zhao et al., 2024) distribute the QA task among multiple LLM-based agents that communicate via message passing, thereby providing greater modularity and a clearer division of labor. By cross-verifying evidence, each agent can potentially identify suspicious partial solutions. However, even these designs often lack a systematic strategy to revert to an earlier, valid state once a global contradiction emerges. Such a reversal capability is non-trivial, as it introduces synchronization complexities among the agents and raises questions about how to detect, prioritize, and handle contradictory or low-confidence information in a large-scale collaborative setting (He et al., 2021).
In this paper, we propose ReAgent, a reversible multi-agent collaborative reasoning framework for multi-hop QA (Figure 1, right). Our method introduces explicit backtracking protocols into a multi-layer architecture that addresses both the granular aspects of local error correction and the broader system-wide consistency checks. ReAgent alleviates error accumulation in forward-only strategies by incorporating fine-grained conflict-detection cues at each step and using a flexible error-correction loop that can revert and iteratively refine intermediate inferences. Our design is inspired by multi-agent systems with reflective capabilities (Bo et al., 2024; Huang et al., 2024), but we go further by defining how local versus global backtracking unfolds, how to manage concurrency issues when parallel agents must revert their states, and how to integrate trust signals derived from each agent’s past reliability (Puerto et al., 2023; Parhizkar et al., 2020). While the addition of reversible reasoning raises legitimate questions about computational overhead and concurrency, our findings show that these challenges can be mitigated by careful coordination at the supervisory level. Experiments on three public multi-hop QA benchmarks demonstrate that ReAgent improves final-answer accuracy while enhancing interpretability, and even outperforms several advanced reasoning models despite it built on a lightweight, non-reasoning foundation.
Figure 2: The overall architecture of the ReAgent. The given question is processed through the Execution Layer, which involves question decomposition, evidence retrieval, verification, and is ultimately integrated to generate the final answer (blue line). The Supervisor Layer and Interaction Layer are responsible for monitoring, regulation, and communication. The ReAgent framework includes both local and global backtracking mechanisms (red boxes), triggered by the Verifier Agent ($A_{V}$) and Supervisor Agent ($A_{S}$), respectively.
Our contributions are threefold: (1) we propose a multi-agent QA framework that supports both local and global backtracking to correct mistakes in situ; (2) we design a hybrid retrieval mechanism that integrates textual and graph-based evidence; and (3) We empirically demonstrate the effectiveness of ReAgent, achieving an average accuracy gain of approximately 6% over the strongest baselines, while improving robustness and transparency compared to forward-only methods.
2 Related Work
Prompting and Iterative Reasoning. Large Language Models (LLMs) can tackle complex tasks using chain-of-thought (CoT) promptingYang et al. (2024a, 2025b), which promotes step-by-step reasoning and improves performance in arithmetic, commonsense, and symbolic tasks (Wei et al., 2022). Accuracy further improves with self-consistency, which samples multiple reasoning paths and selects the majority answer (Wang et al., 2023). To reduce manual prompt design, automatic methods such as APE generate and evaluate prompts via LLMs themselves, treating prompt creation as a search task (Zhou et al., 2023). Iterative frameworks like ReAct combine reasoning with tool use to refine answers (Yao et al., 2023), while Reflexion and Self-Refine add feedback loops where the model critiques and revises its outputs (Shinn et al., 2023; Madaan et al., 2023). Prompting has also extended to multimodal inputs; for instance, Multimodal-CoT integrates visual and textual information into a joint reasoning trace to generate the final answer (Zhang et al., 2023).
Multi-Agent Collaboration, Debate, and Scalability. Recent work investigates multi-agent systems where multiple LLMs collaborate or compete toward shared goals. CAMEL assigns roles (e.g., “user” and “assistant”) to agents that communicate via dialogue to decompose tasks (Li et al., 2023). Role specialization and consensus mechanisms enhance robustness. Debate-based methods push this further, with agents arguing opposing views and a judge—human or model—selecting the best solution (Khan et al., 2024; Xiong et al., 2023). Such adversarial exchanges promote correctness, though dominant agents can bias results without capable judges. As agent counts grow, coordination becomes a bottleneck. Efficient structures like hierarchical controllers and learned protocols are needed to manage scalability and communication overhead (Li et al., 2023; Xiong et al., 2023).
@@ 3 Preliminary
Non-Monotonic Backtracking. Our notion of backtracking is a non-monotonic extension of the typical multi-hop QA process. We introduce a reversible reasoning mechanism that allows both local and global backtracking.
  • Local Correction: An agent can revise its own inference when it detects internal conflict or receives contradictory evidence from other agents.
  • Global Rollback: A supervisor coordinates rollback across agents when inconsistencies span multiple agents, restoring the system to a previously consistent state.
@@ 4 ReAgent: Reversible Multi-Agent Reasoning Architecture
Figure 2 presents the overall architecture of our proposed ReAgent, organized into three layers: 1) Execution Layer, responsible for decomposing the input question $Q$ into multiple sub-questions, retrieving relevant evidence respectively, validating intermediate results, and integrating them to generate the final answer $A_{final}$; 2) Supervisor Layer, responsible for high-level regulation, coordinating conflict resolution and managing global backtracking; and 3) Interaction Layer, responsible for maintaining the concurrency model and communication protocols.
The core of the architecture is a hierarchical backtracking mechanism, consisting of local backtracking, which resolves internal contradictions within each agent, and global backtracking, which handles contradictions spanning multiple agents. To support this, ReAgent maintains knowledge sets at each time step $t$. Specifically, given a set of agents $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$, each agent $A_{i}$ holds a local knowledge set $\Phi_{i}^{t}$, containing its proposed assertions, retrieved evidence, or intermediate inferences, while the system maintains a global knowledge set $\Phi^{t}=\bigcup_{i=1}^{n}\Phi_{i}^{t}$, representing the set of global statements under consideration at time $t$. The design explicitly supports non-monotonic updates: newly introduced statements can be revoked if they lead to logical conflicts or are superseded by contradictory evidence. The specific prompts of each kind of agent are provided in Appendix A.
@@ 4.1 Execution Layer Agents
The Execution Layer hosts four types of agents that address fundamental QA sub-tasks, each maintaining its local knowledge $\Phi_{i}$.
Question-Decomposer Agent ($A_{Q}$). Given a complex input question $Q$, this agent breaks it into a set of sub-questions for subsequent retrieval and verification: $A_{Q}:Q\longmapsto\{q_{1},\dots,q_{m}\}$. The decomposition is broadcast to other base-layer agents.
Retriever Agent ($A_{R}$). Upon receiving a sub-question $q_{i}$, $A_{R}$ issues parallel sparse and dense queries over the corpus, merges the hits with reciprocal-rank fusion, and retains the top-$M$ passages as the evidence set $E={e_{1},\dots,e_{M}}$. If backtracking invalidates any $e_{j}\in E$, only the associated $q_{i}$ is re-queried, leaving the rest unchanged.
Verifier Agent ($A_{V}$). Performs local consistency checks. Given a new evidence set $E$, it verifies coherence with its current knowledge $\Phi_{V}^{t}$. If conflicts arise, it triggers local backtracking to revert to a consistent state or signals higher-level agents for broader resolution. Concretely, $A_{V}$ may invoke $\mathrm{BacktrackLocal}\bigl{(}A_{V},r\bigr{)}$ if it detects inconsistencies in assertions introduced after node $r$.
Answer-Assembler Agent ($A_{A}$). Gathers partial answers (and verified evidence) to synthesize a final answer. Given the local inferences from $A_{Q}$, $A_{R}$, and $A_{V}$, the agent combines them to generate the final answer $A_{final}$, represented as:
$A_{final}=A_{A}(\Phi_{Q},\Phi_{R},\Phi_{V}),$
where $\Phi_{Q}$, $\Phi_{R}$ and $\Phi_{V}$ denotes the knowledge set from the respective agents. Any detected contradiction spanning multiple agents triggers escalation to the supervisory layer.
@@ 4.2 Supervisory Layer Agents
The supervisory layer oversees system-wide strategies, especially when contradictory goals or inconsistent states appear across multiple agents that cannot be resolved by a single agent itself.
Controller Agent ($A_{C}$). Regulates high-level strategies by monitoring game-theoretic signals or meta-rules. If a certain rule is deemed suboptimal or unsafe, $A_{C}$ can override it or enforce mode switches. For instance, it may issue a $\textsf{challenge}(\varphi)$ to examine a crucial assertion $\varphi$ from multiple agents’ perspectives or override local decisions if they jeopardize overall consistency.
Supervisor Agent ($A_{S}$). Coordinates multi-agent conflicts spanning multiple knowledge sets or critical shared assumptions. If a conflict persists after local backtracking and escalation from $A_{A}$, the agent determines whether partial or holistic rollback is required. Specifically, when a system-wide contradiction is found, $A_{S}$ identifies a minimal conflict set $\Psi$ such that $\mathrm{SAT}(\Psi)=\textsf{false}$ but $\mathrm{SAT}(\Psi^{\prime})=\textsf{true}$ for any proper subset $\Psi^{\prime}\subset\Psi$. The supervisor then triggers a backtracking operation that eliminates or modifies $\Psi$.
@@ 4.3 Interaction Layer
The Interaction Layer is responsible for storing knowledge sets from each interaction round, as well as maintaining the concurrency model and communication protocols.
Persistent Log. Stores the local knowledge sets $\Phi_{i}$ and global knowledge set $\Phi$ from all interaction rounds to support backtracking and serve as a historical evidence repository.
Temporal Tracker. Records the chronological sequence of messages and actions, enabling agents to reference previous steps accurately. Specifically, the system can use temporal operators such as $\square\varphi$ to denote that $\varphi$ must hold in every future state, and $\lozenge\varphi$ to indicate that a proposition $\varphi$ might become true at some future point. These temporal constraints assist in specifying persistent axioms or potential triggers for backtracking conditions.
Messaging Channel. Achieves communication across agents for exchanging updates, signaling conflicts, and broadcasting intermediate conclusions. Specifically, atomic events $\mathtt{msg}(\varphi)$ are typed as assert, inform, reject, or challenge. These updates can occur simultaneously, and a composite event model merges parallel messages. Furthermore, a shared concurrency mechanism processes simultaneous actions, ensuring that conflicts arising from concurrent assertions are escalated to the supervisory layer.